=================================================================== Tohban report: 2009-02-17 to 2009-02-23 (Larry Kepko) Produced at ? Checked data from: 2009-02-15/00:00 UT to 2009-02-21/24:00 UT Last effective data dump on TH-B ? =================================================================== Memory Status A: 02/15-02/21 ok B: 02/15-02/19 12:00 ok filled 02/19 12:00 to 02/20 07:00 C: 02/15-02/21 ok D: 02/15-02/21 ok E: 02/15-02/21 ok FGM: A: 02/15-02/21 full coverage B: 02/15-02/19 full coverage no data after 02/20 03:00 C: 02/15-02/21 18:45 full coverage D: 02/15-02/21 15:00 full coverage E: 02/15-02/21 16:30 full coverage ESA: A: 02/15-02/21 full coverage B: 02/15-02/17 09:30 full coverage 02/17 09:30-02/18 20:30 no data 02/18 20:30 02/19 12:00 full coverage C: 02/15-02/18 01:00 full coverage 02/18 01:00-02/18 20:30 no data 02/18 20:30-02/20 06:30 full coverage 02/20 06:30-02/20 20:30 no data 02/20 20:30-02/21 22:30 full coverage D: 02/15-02/21 19:00 full coverage E: 02/15-02/21 20:30 full coverage except 02/20 20:30-02/20 24:00 SST: A: 02/15-02/21 full coverage B: 02/15-02/17 09:30 full coverage 02/17 09:30-02/18 20:30 no data 02/18 20:30 02/19 12:00 full coverage C: 02/15-02/18 01:00 full coverage 02/18 01:00-02/18 20:30 no data 02/18 20:30-02/20 06:30 full coverage 02/20 06:30-02/20 20:30 no data 02/20 20:30-02/21 22:30 full coverage D: 02/15-02/21 19:00 full coverage E: 02/15-02/21 20:30 full coverage except 02/20 20:30-02/20 24:00 EFI filter bank: A: 02/15-02/21 full coverage B: 02/15-02/19 12:00 full coverage C: 02/15-02/21 22:30 full coverage D: 02/15-02/21 19:00 full coverage E: 02/15-02/21 20:30 full coverage SCM filter bank: A: 02/15-02/21 full coverage B: 02/15-02/19 12:00 full coverage C: 02/15-02/21 22:30 full coverage D: 02/15-02/21 19:00 full coverage E: 02/15-02/21 20:30 full coverage #--------------------------------------------------------------------------------- # 2009/046 - 2009/047 Sunday, February 15 #--------------------------------------------------------------------------------- BGS_11M THEMIS_C -9.400 2009/047 02:39:20 2009/047 03:21:00 TR10 # Data dump [ME] - change data rate to TR9 at 03:01 #--------------------------------------------------------------------------------- # 2009/047 - 2009/048 Monday, February 16 Week 8 #--------------------------------------------------------------------------------- #--------------------------------------------------------------------------------- # 2009/048 - 2009/049 Tuesday, February 17 #--------------------------------------------------------------------------------- MILA_9M2 THEMIS_B 2.980 2009/048 23:46:00 2009/049 00:16:00 TR2 # SOH check and track [DP] - upload ETC table version 8 to EEPROM and IDPU script set 0009 to RAM - ICCR_DOP78 & ICCR_DOP79 USNAU_13M THEMIS_C 9.400 2009/049 00:44:00 2009/049 01:14:00 TR10 # Data dump [DP] - upload ETC table version 8 to EEPROM and IDPU script set 0009 to RAM - ICCR_DOP78 & ICCR_DOP79 #--------------------------------------------------------------------------------- # 2009/049 - 2009/050 Wednesday, February 18 #--------------------------------------------------------------------------------- #--------------------------------------------------------------------------------- # 2009/050 - 2009/051 Thursday, February 19 #--------------------------------------------------------------------------------- AGO_9M THEMIS_A 9.100 2009/050 22:09:00 2009/050 22:23:00 TR10 # Data dump - negative acquisition due to ITOS hanging MILA_9M2 THEMIS_A -9.100 2009/050 22:50:00 2009/050 23:08:00 TR9 # Data dump [JT] - blind acqusition - all data retrieved except 1 burst AGO_9M THEMIS_B 9.100 2009/051 06:45:00 2009/051 07:15:00 TR10 # Data dump [JM] - lost 1% of VC3 frames due to link problems #--------------------------------------------------------------------------------- # 2009/051 - 2009/052 Friday, February 20 #--------------------------------------------------------------------------------- BGS_11M THEMIS_D 9.300 2009/051 19:58:55 2009/051 20:23:55 TR10 # Data dump [ML] - upload and enable new RTS 18, 19, 20 and 29 USNHI_13M THEMIS_E 2.550 2009/051 20:35:00 2009/051 20:55:00 TR2 # SOH check and track [ML] - LRO proficiency support - upload and enable new RTS 18, 19, 20 and 29 #--------------------------------------------------------------------------------- # 2009/052 - 2009/053 Saturday, February 21 #--------------------------------------------------------------------------------- 02/15 - Plasmasheet oscillations 05:00-10:00 02/15 - Tailward flow TH-B, 03:10 02/17 - TH-A/C flows, 03:45 02/21 - Multiple flow bursts TH-C, dipolarizations inner probes